# Contextual modulation of language comprehension in a dynamic neural model of lexical meaning

### v1

@@ Abstract
We propose and computationally implement a dynamic neural model of lexical meaning, and experimentally test its behavioral predictions. We demonstrate the architecture and behavior of the model using as a test case the English lexical item have, focusing on its polysemous use. In the model, have maps to a semantic space defined by two continuous conceptual dimensions, connectedness and control asymmetry, previously proposed to parameterize the conceptual system for language. The mapping is modeled as coupling between a neural node representing the lexical item and neural fields representing the conceptual dimensions. While lexical knowledge is modeled as a stable coupling pattern, real-time lexical meaning retrieval is modeled as the motion of neural activation patterns between metastable states corresponding to semantic interpretations or readings. Model simulations capture two previously reported empirical observations: (1) contextual modulation of lexical semantic interpretation, and (2) individual variation in the magnitude of this modulation. Simulations also generate a novel prediction that the by-trial relationship between sentence reading time and acceptability should be contextually modulated. An experiment combining self-paced reading and acceptability judgments replicates previous results and confirms the new model prediction. Altogether, results support a novel perspective on lexical polysemy: that the many related meanings of a word are metastable neural activation states that arise from the nonlinear dynamics of neural populations governing interpretation on continuous semantic dimensions.
@@
lexical semantics , dynamic field theory , language comprehension[inst1]organization=Linguistics Department, Yale University,addressline=370 Temple St., city=New Haven, state=Connecticut, postcode=06511, country=USA
We propose a model of lexical meaning in the framework of Dynamic Field Theory (DFT)
@@
We argue that polysemy arises from metastability in neural dynamics on continuous dimensions
@@ 1 Introduction
In language comprehension, linguistic forms evoke interpretations of meaning. A basic unit of linguistic form is the lexical item, the systematic relationship between phonetic/phonological, morphological, syntactic, and semantic information (Anderson, 1992; Jackendoff, 1975). It is a well-known observation that the specific interpretation or reading evoked by a lexical item can vary depending on context. Polysemy refers to the situation whereby a lexical item offers more than one interpretation yet those possibilities are conceptually related. For example, the lexical item book can be described as polysemous since its possible interpretations include (at least) both a physical object (e.g., heavy book) and a collection of information represented by an object (e.g., enjoyable book) (e.g., Brugman, 1988; Deane, 1988; Lakoff, 1990; Pustejovsky, 1995; Vicente, 2018).111Polysemy contrasts with homophony, the situation where the possible interpretations appear semantically unrelated, e.g., river bank vs. savings bank. Here we investigate the conceptual basis of the constrained semantic variability observed in lexical polysemy. To this end, we examine a possible source of this variability, propose a neuro-computational implementation, and investigate the behavioral consequences during sentence comprehension. Under our proposal, the many-to-one relation between meaning and form which is salient in cases of lexical polysemy can be seen as the norm rather than the exception, with differences primarily in the magnitude of variability and the relatedness of the possible interpretations.
@@ 1.1 English ‘have’
. The linguistics professor has a motorcycle. (alienable possession)
@@
When subjects are presented with less frequent argument combinations, as in 1.1 below, two types of responses are reported: (a) the inanimate subject “the oak tree” leads to an inalienable possession interpretation, but this conflicts with the fact that the object “the motorcycle” does not plausibly enter into a part-whole relation with “the oak tree”, leading to a decrease in acceptability; (b) the referent of the subject is anthropomorphized, e.g., “the person dressed as a tree…”, in order to maintain an alienable possession interpretation (Zhang et al., 2018; Zhang, 2021). These facts are consistent with a bias for have to evoke a possession interpretation.
. #The oak tree has a motorcycle.222“#” signifies that the sentence, while grammatical, is judged less felicitous without additional context.
But the polysemy reportoire of have is larger than these two readings. have can also evoke an interpretation of coincidental location. This would render 1.1 above with an interpretation that the motorcycle is in coincidental spatial proximity to (e.g., next to, under, below, above) the oak tree, a situation that we will refer to as adjacency (e.g., Myler, 2016; Zhang et al., 2022). The tendency for have to evoke a possession reading is very strong, but not indefeasible. One standard way to bring out an adjacency reading from have is to add a prepositional phrase which makes explicit the spatial relation, as in 1.1. Not surprisingly, presence of the modifier “next to it” improves acceptability ratings relative to sentences like 1.1 (Zhang et al., 2018, 2022).
@@
This said, presence of an explicit modifier is not the only way to make an adjacency reading salient. A bias for an adjacency interpretation can also be induced by the preceding context. Specifically, when a preceding sentence evokes an adjacency interpretation, as in 1.1, then sentences like 1.1 are more likely to trigger an adjacency interpretation as well as receive higher acceptability ratings (Zhang et al., 2018, 2022), as seen in 1.1.
@@
This linguistic and behavioral pattern suggests a unified polysemy-based analysis of English have involving readings raging from a purely coincidental spatial relation to an alienable possession relation to an inalienable possession relation. Crucially, these readings are not disconnected. The distinction between adjacency, alienable possession, and inalienable possession is a matter of degree, not category, not only between these three reading types but also within them. In the case of coincidentality in spatial location, for example, the situation conveyed by “the tree has a motorcycle” is more coincidental than “the tree has a bench”, and that, in turn, is more coincidental than “the tree has a swing”, or “the tree has a nest”. So, with each situation we move towards less coincidentality and, interestingly, towards greater asymmetry between arguments such that the first argument is more and more perceived as possessing the second.
Similar gradability is observable within the alienable and inalienable possession spaces. For alienable possession, “the woman has a car” can refer to a variety of related situations distinguishable by degree of ownership: the woman has a car because she stole it, because she borrowed it, rented it, bought it. Only in the last instance would she “own” the car, yet in all instances the relation between the woman and the car—one of asymmetry in control—renders it one of possession. Finally, for the inalienable possession space, the possibilities range from the peripheral “the woman has hair” to the more inalienable “the woman has a liver”, to the completely inalienable “the woman has a body”. Notably, while there is no coincidentality in these examples, making it more like the alienable possession space, the control relation between arguments—e.g., woman to her hair, woman to her liver, woman to her body—has become less and less asymmetrical, similar to the relation in the coincidental location space (Koch, 2012; Piñango, 2023; Zhang, 2021).
We conclude then that the lexical meaning of have, while constrained, is variable along continuous dimensions. As a result, such meaning is best characterized as a space within which specific readings can obtain. This conclusion in turn raises the question of what the properties of such a continuous space are that give rise to the constrained variability in meaning that have exhibits. We address this question below.
@@ 1.2 Continuous conceptual space for lexical meaning
It has been proposed that the readings of have are supported by a conceptual space organized in terms of two continuous cognitive dimensions or metrics: control asymmetry and connectedness (Gärdenfors and Warglien, 2012; Piñango, 2019, 2023; Zhang, 2021). Control asymmetry measures the degree to which two individuals in a situation differ in how much one controls the other. Control asymmetry underpins perception of causality and causal chains. In doing so, it gives rise to intuitions about intentionality and agency (e.g., Carey, 2009; Croft, 2012; Talmy, 1988). It has also been shown to constrain language development (e.g., Klein and Perdue, 1992). An evaluation of high control asymmetry between two participants means that one of the participants can be construed as a controller in the situation, that is, an intentional agent, and not the other way around. A low control asymmetry evaluation means that neither of the participants is construable as controller. In the case of have, ultimate evaluation of control asymmetry depends on the context, including the properties of the participants.
Connectedness measures the degree to which two individuals in a situation are functionally or structurally related to one another. It is from connectedness that intuitions emerge about coherence in the world, e.g., object individuation (e.g., Carey, 2009; Gopnik et al., 2004; Krøjgaard, 2004). Relations of connectedness are built along independently motivated conceptual dimensions—spatial, temporal, informational, or functional (Piñango and Deo, 2016). In the case of have, ultimate evaluation of connectedness again depends on the context and the perceived properties of the argument referents.
In Figure 1, three possible interpretations of have are plotted in the space spanned by the two dimensions of connectedness (horizontal axis) and control asymmetry (vertical axis). A coincidental adjacency interpretation, as in 1.1, corresponds to low connectedness and low control asymmetry, and an alienable possession interpretation, as in 1.1, corresponds to higher connectedness and higher control asymmetry. An inalienable possession interpretation, as in 1.2, corresponds to high connectedness and low control asymmetry.
@@
As seen in Figure 2, when more possible interpretations of have are plotted (including, in addition to the examples given so far, at least containment, control, and kinship), they tend to fall on an inverted “U” shape. This pattern indicates dependencies between the two dimensions which constrain object perception and its linguistic encoding. We return to this point in the context of our model in Section 2.
1.3 Further empirical evidence for continuous semantic dimensions
In addition to the robust observation of gradability between readings of have, empirical support for a model of lexical meaning based on interpretable, continuous dimensions comes from two other sources: (a) trajectories of meaning change over time, and (b) real-time sentence comprehension effects. Regarding (a), over an approximately 200-year period, the postposition kade in the Indo-Aryan language Marathi has shifted gradually from primarily an adjacency reading, to primarily an alienable possession reading, to primarily an inalienable possession reading (Deo, 2015; Zhang, 2021). This trajectory essentially traces the inverted “U” shape in Figure 2 from left to right. This supports the notion that the potential readings of have are organized along these continuous dimensions, rather than forming a countable set of discrete meanings.
@@ 1.3 Further empirical evidence for continuous semantic dimensions
Regarding (b), evidence comes from acceptability judgments, self-paced reading, and electroencephalography (EEG) experiments investigating the availability of the coincidental adjacency reading associated with have as a result of preceding context. Zhang et al. (2018) administered an acceptability judgment task with target sentences consisting of have and two inanimate arguments, as in 1.1. Each target sentence was preceded by a context sentence, also consisting of have and two inanimate arguments. The crucial manipulation was the meaning of the context sentence: adjacency, as in 1.1, or inalienable possession, as in 1.3.
@@
When preceded by a context sentence designed to evoke an adjacency interpretation, the target sentence received higher acceptability ratings, relative to when the same sentence was preceded by a sentence with a possession meaning. This improvement in acceptability was attributed to a gradient shift in interpretation towards adjacency due to the influence of the context sentence.
@@
Using a similar stimulus set as in the acceptability judgment task, Zhang et al. (2018) observed decreased reading times for target sentences preceded by an adjacency context sentence, relative to when they were preceded by a possession context sentence. Moreover, EEG recordings indicated an N400 event-related potential (ERP) in the possession condition relative to the adjacency condition (Zhang et al., 2018). These results are consistent with a decreased neurocognitive processing load in the adjacency condition, due to contextual facilitation of an adjacency interpretation of the target sentence.
@@
Finally, the magnitude of contextual modulation of have interpretation correlates at an individual level with an independent measure of context sensitivity, i.e., the Autism-Spectrum Quotient (AQ: Baron-Cohen et al., 2001). Individuals with higher AQ scores (decreased context sensitivity) exhibit a decreased difference in acceptability between conditions (Zhang et al., 2022). The influence of context on interpretation is a gradient rather than categorical phenomenon. Some of the variation in the magnitude of contextual influence can be explained by the AQ.
Taken together, observations from historical change, acceptability judgments, self-paced reading, EEG, and individual variation all point to gradience in the cognitive representation of lexical meaning. Such gradience is captured by a parameterized conceptual space such as the one schematized in Figures 1 and 2. The question that arises now is what are the neurocognitive properties of such a system. To investigate this question we turn to Dynamic Field Theory (DFT: Schöner et al., 2016), a formal framework for understanding the neural activation dynamics underlying continuous cognitive dimensions.

### v2

@@ Abstract
We computationally implement and experimentally test the behavioral predictions of a dynamic neural model of lexical meaning in the framework of Dynamic Field Theory. We demonstrate the architecture and behavior of the model using as a test case the English lexical item have, focusing on its polysemous use. In the model, have maps to a semantic space defined by two independently motivated continuous conceptual dimensions, connectedness and control asymmetry. The mapping is modeled as coupling between a neural node representing the lexical item and neural fields representing the conceptual dimensions. While lexical knowledge is modeled as a stable coupling pattern, real-time lexical meaning retrieval is modeled as the motion of neural activation patterns between transiently stable states corresponding to semantic interpretations or readings. Model simulations capture two previously reported empirical observations: (1) contextual modulation of lexical semantic interpretation, and (2) individual variation in the magnitude of this modulation. Simulations also generate a novel prediction that the by-trial relationship between sentence reading time and acceptability should be contextually modulated. An experiment combining self-paced reading and acceptability judgments replicates previous results and partially bears out the model’s novel prediction. Altogether, results support a novel perspective on lexical polysemy: that the many related meanings of a word are not categorically distinct representations; rather, they are transiently stable neural activation states that arise from the nonlinear dynamics of neural populations governing interpretation on continuous semantic dimensions.
@@
lexical semantics , polysemy , dynamical systems , dynamic field theory , language comprehensionWe propose a model of lexical meaning in the framework of Dynamic Field Theory (DFT)
@@
We argue that polysemy arises from nonlinear neural dynamics on continuous dimensions
@@ 1 Introduction
In language comprehension, linguistic forms evoke interpretations of meaning. A basic linguistic unit is the lexical item, a systematic relation between phonetic/phonological, morphological, syntactic, and semantic information (e.g., Anderson, 1992; Chomsky, 1965; Jackendoff, 1975). It is well-attested that the specific interpretation or reading evoked by a lexical item can vary depending on context. Polysemy refers to the situation whereby a lexical item offers more than one interpretation yet those possibilities are conceptually related. For example, the lexical item book can be described as polysemous since its possible readings include (at least) both a physical object (e.g., heavy book), and information represented by the object (e.g., enjoyable book) (e.g., Brugman, 1988; Deane, 1988; Lakoff, 1990; Pustejovsky, 1995; Vicente, 2018).111Polysemy contrasts with homophony, the situation where the possible readings appear semantically unrelated, e.g., river bank vs. savings bank. Here we investigate the semantic basis of the constrained variability observed in lexical polysemy. To this end, we examine a possible source of this variability, propose a neuro-computational implementation in the framework of Dynamic Field Theory (Schöner et al., 2016), and investigate its behavioral predictions during sentence comprehension. Under our proposal, the many-to-one relation between meaning and form which is salient in cases of lexical polysemy represents the norm rather than the exception, with differences primarily in the magnitude of variability and the relatedness of the possible readings.222This work is situated in an analytical approach to lexical polysemy whereby meaning variability depends on contextual satisfaction of discrete meaning representations that are, as it were, sequestered in the mental lexicon, and connected to a unified bundle of morphosyntactic and morphophonological properties (e.g., Deo and Pinango, 2011; Jackendoff, 1975, 1997; Piñango and Deo, 2016; Pustejovsky, 1995). Our proposal here shares with that traditional approach the assumption that the meaning generation process involved in polysemy is grounded in the combinatorial properties of the semantic system. It departs from it by placing the burden of meaning generation, including polysemy, on the interaction between the conceptual context of a lexical item and a continuous meaning space. The result of this interaction is what gives the language user the experience of meaning discreteness: what we refer to as a lexical meaning, which is nonetheless never isolated from the larger continuous meaning space. On this view, meaning discreteness is not an input to semantic composition, contrary to what is traditionally assumed, but an outcome of the compositional process itself (Piñango, 2023). To our knowledge, this is the first formal neuro-computational exploration of this idea.
@@ 1.1 English ‘have’
. The oak tree has colorful decorations. (alienable possession)
@@
When people are presented with less frequent argument combinations, as in 1.1 below, two types of responses are reported: (a) the inanimate subject “the oak tree” leads to an inalienable possession reading, which conflicts with the fact that the object “the motorcycle” does not plausibly enter into a part-whole relation with “the oak tree”, leading to a decrease in acceptability; or (b) the referent of the subject is anthropomorphized, e.g., “the person dressed as a tree…”, which maintains an alienable possession interpretation (Zhang et al., 2018; Zhang, 2021). The facts presented so far are consistent with the hypothesis that have primarily evokes possession readings.
. #The oak tree has a motorcycle.333“#” signifies that the sentence, while grammatical, is judged less felicitous without additional context.
But the polysemy reportoire of have is larger than these two readings. have can also evoke an interpretation of coincidental location. This would render 1.1 above with an interpretation that the motorcycle is in coincidental spatial proximity to the oak tree (e.g., next to, under, below, above), an interpretation that we will refer to as adjacency (e.g., Myler, 2016; Zhang et al., 2022). The tendency for have to evoke a possession reading is very strong, but not indefeasible. One standard way to bring out an adjacency reading from have is to add a prepositional phrase which makes explicit the spatial relation, as in 1.1. Not surprisingly, inclusion of the locative modifier “next to it” is reported to improve acceptability ratings relative to sentences like 1.1 (Zhang et al., 2018, 2022).
@@
Inclusion of an explicit locative modifier is not the only way to make an adjacency reading salient. A bias for an adjacency interpretation can also be induced by the preceding context. Specifically, when a preceding sentence evokes an adjacency reading, as in 1.1, then sentences like 1.1 are more likely to evoke an adjacency reading as well as receive higher acceptability ratings (Zhang et al., 2018, 2022), as seen in 1.1.
@@
This linguistic and behavioral pattern suggests a unified polysemy-based analysis of English have involving readings ranging from a purely coincidental spatial relation to an alienable possession relation to an inalienable possession relation (Zhang, 2021). Crucially, these readings are not disconnected. Evidence suggests that the distinction between adjacency, alienable possession, and inalienable possession is a matter of degree, not category, not only between these three reading types but also within them (e.g., Deo, 2015; Zhang, 2021; Zhang et al., 2022). In the case of coincidentality in spatial location, for example, the situation conveyed by “the tree has a motorcycle” is more coincidental than “the tree has a bench”, and that, in turn, is more coincidental than “the tree has a swing”, or “the tree has a nest”. So, with each situation we move towards less coincidentality and, interestingly, towards greater asymmetry between arguments such that the first argument is more and more perceived as “possessing” the second.
Similar gradability is observable within the alienable and inalienable possession spaces. For alienable possession, “the woman has a car” can refer to a variety of related situations distinguishable by degree of ownership: the woman has a car because she stole it, because she borrowed it, rented it, bought it. Only in the last instance would she “own” the car, yet in all instances the relation between the woman and the car—one of asymmetry in control—renders it one of possession. Finally, for the inalienable possession space, the possibilities range from the peripherally inalienable “the woman has hair” to the more inalienable “the woman has a liver”, to the completely inalienable “the woman has a body”. Notably, while there is no coincidentality in these examples, making it more like the alienable possession space, the control relation between arguments—e.g., woman to her hair, woman to her liver, woman to her body—has become less and less asymmetrical, mirroring the relation in the coincidental location space (Koch, 2012; Piñango, 2023; Zhang, 2021).
We conclude then that the lexical meaning of have, while constrained, gives rise to gradable readings. As a result, such meaning is best characterized as a continuous space within which specific readings can obtain. This conclusion in turn raises the question of what the properties of such a continuous space are. We address this question below.
@@ 1.2 Continuous conceptual space for lexical meaning
It has been proposed that the readings of have are supported by a conceptual space organized in terms of two continuous cognitive dimensions or metrics: control asymmetry and connectedness (Piñango, 2019, 2023; Zhang, 2021, see also Gärdenfors (2000, 2014) for the notion of continuous conceptual dimensions and their relation to linguistic semantics).444In linguistics, the notion that continuous dimensions underlie apparently categorical distinctions has also been invoked in the domain of phonology (e.g., Browman and Goldstein, 1989; Pierrehumbert et al., 2011; Stevens, 1989). At the level of neural encoding, it is uncontroversial that dimensions underlying perception (e.g., Dyballa et al., 2024; Hubel and Wiesel, 1959; Leonard et al., 2024) and movement (e.g., Bouchard et al., 2013; Chartier et al., 2018; Georgopoulos et al., 1986) are continuous. Control asymmetry measures the degree to which two individuals in a situation differ in how much one controls the other. Control asymmetry underpins perception of causality and causal chains. In doing so, it gives rise to intuitions about intentionality and agency (e.g., Carey, 2009; Croft, 2012; Talmy, 1988). It has also been shown to constrain language creation in pidgins and creoles (e.g., Klein and Perdue, 1992). An evaluation of high control asymmetry between two participants means that one of the participants can be construed as a controller in the situation, and not the other way around. A low control asymmetry evaluation means that neither of the participants is construable as controller.555In the example sentences so far, as well as in the stimulus sentences in the experiment described below, both arguments of have denote inanimate entities. Even though “control” normally evokes animacy, it is not limited to it. Control asymmetry exists between inanimate entities whenever the state of one can be seen as depending on the other, and not vice versa; or when one can exert influence on the other, and not vice versa. In 1.1, for example, the tree exerts greater influence on the decorations than vice versa, since the state of the decorations (in particular, the spatial position) depends more on the tree (since they are hanging on it) than vice versa, as the state of the tree would remain unchanged if the decorations were removed.
Connectedness measures the degree to which two individuals in a situation are functionally or structurally related to one another. It is from connectedness that intuitions emerge about coherence in the world, e.g., object individuation and part-whole relations (e.g., Carey, 2009; Gopnik et al., 2004; Krøjgaard, 2004). Relations of connectedness are built along independently motivated conceptual dimensions—spatial, temporal, informational, or functional (Piñango and Deo, 2016).
In Figure 1, three possible readings of have are plotted in the space spanned by the two dimensions of connectedness (horizontal axis) and control asymmetry (vertical axis). A coincidental adjacency reading, as in 1.1, corresponds to low connectedness and low control asymmetry, and an alienable possession reading, as in 1.1, corresponds to higher connectedness and higher control asymmetry. An inalienable possession reading, as in 1.2, corresponds to high connectedness and low control asymmetry.
@@
As seen in Figure 2, when more possible readings of have are plotted (including, in addition to the examples given so far, at least containment, control, and kinship), they tend to fall on a downward parabola. This pattern indicates dependencies between the two dimensions which constrain object perception and its linguistic encoding. We return to this point in the context of our model in Section 2.
In addition to the robust observation of gradability between readings of have, empirical support for a model of lexical meaning based on interpretable, continuous dimensions also comes from trajectories of meaning change over time. Over an approximately 200-year period, the postposition kade in the Indo-Aryan language Marathi has shifted gradually from primarily an adjacency reading, to primarily an alienable possession reading, to primarily an inalienable possession reading (Deo, 2015; Zhang, 2021). This trajectory essentially traces the downward parabola shape in Figure 2 from left to right. This supports the notion that the potential readings of have are organized along these continuous dimensions, rather than forming a countable set of discrete meanings.666An anonymous reviewer notes that the adjacency and possession readings that we have focused on here hardly exhaust the polysemic repertoire of have. Cases like they have a miscarriage/a conniption/a seizure or cases like have mercy/pity/compassion (in the imperative form), or they have a banana/a bite/dinner are also available readings, yet at first glance they don’t fall within either of those groupings. Full discussion of these cases is outside of the scope of the paper. However, we want to provide a sense of how they are captured within the space. The first three cases represent inalienable possession readings (high connectedness, low control asymmetry). have a miscarriage is a variant in that the eventuality of miscarriage demands a shift in connectedness (from high to mid) implementing the process of alienation between the mother and the fetus. In the second set, mercy, pity and compassion are emotional manifestations of a mind, indicating high connectedness and normally low control asymmetry. Yet, the use in the imperative construction conveys the possibility of control asymmetry at some level. Finally, the consumption readings (have a banana/bite/dinner) involve a process (the act of ingesting) that results in a state: the point after ingestion has taken place, resulting in the theme changing location. As in the case of miscarriage, these readings demand a shift, a move from mid to higher connectedness and to lower control asymmetry (closer to the part-whole/inalienable possession space), revealing the change in relation between the two participants that ingestion creates. This is because whatever is ingested becomes part of the organism that ingested it and ultimately can have the power to “control” such organism. This is evident in cases like, e.g., have the hemlock where the controllee once ingested becomes the controller of sorts of the host organism.777Some aspectual uses of have appear to be outside the uses alluded to here. Indeed, the auxiliary uses represent a further grammaticalization of have brought about by morphosyntactic reanalysis during Old English of the object and the past participle; a reanalysis which was made possible by the freer word order of the language at that time (e.g., Michaelis, 1993; Pancheva, 2003). Specifically, the English have-perfect appears to have emerged from a reanalysis of the complex transitive use, indicating possession—as in “I [havev [[a houseobj] [brokenadj]]”—to a transitive use where the possession reading is no longer available, as in “I [have [broken [a house]]]”. This left have to take on an additional semantically distinct use, i.e., end sub-interval of the event, thus creating a novel lexical item dedicated to this semantic use but only when appearing in this “split predication” configuration (ibid.).
1.3 Real-time comprehension of have
@@
A range of real-time comprehension evidence is consistent with the hypothesis that the lexical meaning of have is defined on these two continuous dimensions. Such evidence comes from acceptability judgments, self-paced reading, and electroencephalography (EEG) experiments investigating the availability of the coincidental adjacency reading associated with have as a result of preceding context. Zhang et al. (2018) administered an acceptability judgment task with target sentences consisting of have and two unrelated, inanimate arguments, as in 1.1. Each target sentence was preceded by a context sentence, also consisting of have and two inanimate arguments. The crucial manipulation was the meaning of the context sentence: adjacency, as in 1.1, or inalienable possession, as in 1.3.
@@
When preceded by a context sentence designed to evoke an adjacency reading, the target sentence received higher acceptability ratings, relative to when the same target sentence was preceded by a context sentence designed to evoke a possession reading. This improvement in acceptability was attributed to a gradient shift in interpretation towards adjacency due to the influence of the context sentence.
@@
Using a similar stimulus set as in the acceptability judgment task, Zhang et al. (2018) observed decreased reading times for target sentences preceded by an adjacency context sentence, relative to when they were preceded by a possession context sentence. Moreover, EEG recordings indicated an N400 event-related potential (ERP) in the possession condition relative to the adjacency condition (Zhang et al., 2018). These results are consistent with a decreased neurocognitive processing load in the adjacency condition, due to contextual facilitation of an adjacency reading of the target sentence.
@@
Finally, the magnitude of contextual modulation of have interpretation correlates at an individual level with an independent measure of context sensitivity, i.e., the Autism-Spectrum Quotient (AQ: Baron-Cohen et al., 2001). Individuals with higher AQ scores (decreased context sensitivity) exhibit a decreased difference in acceptability between conditions (Zhang et al., 2022). These results suggest that the influence of context on interpretation is a gradient rather than categorical phenomenon, and some of the variation in the magnitude of contextual influence can be explained by the AQ. However, these results do not necessarily rule out a model based on discrete meaning representations. We discuss this alternative possibility below.
1.4 A discrete alternative
One could propose that the meaning of have is represented as a countable set of discrete meanings, including, at least, possession and adjacency. Preceding context would bias the comprehender towards one of these discrete options, and when the target sentence violates the comprehender’s expectation, surprisal would result in lowered acceptability, slower reading, and modulation of electrophysiological responses (e.g., Frazier, 1987; Levy, 2008). Gradience in such measurements could result from randomly distributed noise around underlyingly discrete responses types.
Such a proposal is particularly appealing when only two possible readings are considered, as in the experiments described above, as well as in the experiment reported below. However, the broader range of empirical evidence reviewed above supports the continuous account: the synchronic range of have interpretation, the diachronic trajectory of change in the meaning of Marathi kade, and evidence for continuous cognitive dimensions in other domains. The continuous representational account also has two theoretical advantages over a discrete representational account: parsimony and explanatory power. Under a discrete representational account, a large number of representations would have to be posited: one for each reading. The continuous account is more parsimonious because it includes fewer representations: only two (continuous) dimensions. It is also more explanatory because it explicitly relates the different readings to each other, via direction and distance in the continuous space. Ultimately, extending our approach to the larger variety of possible interpretations of have–—as well as other cases of polysemy—–will be necessary to support the hypothesis of continuous representations for lexical meaning. A main contribution of this paper is to offer a neural processing model of lexical meaning on continuous dimensions, which is a necessary step towards extending to other interpretations and other lexical items. To situate our neural processing model, we turn now to Dynamic Field Theory (DFT: Schöner et al., 2016), a formal framework for understanding the neural activation dynamics underlying continuous cognitive dimensions.